OpenStudy (anonymous):

A bag contains a mixture of almonds, walnuts, and cashews. There is an equal quantity of each nut. If a nut is chosen randomly out of the bag, what is the probability that the nut is either an almond, a walnut, or a cashew? Round answer to the nearest thousandth. A.0.333 B.0.667 C.0.750 D.1

OpenStudy (kropot72):

The required probability is the sum of the probabilities of each possible outcome, the reason being that the outcomes are mutually exclusive. \[P(one\ of\ 3\ types)=\frac{1}{3}+\frac{1}{3}+\frac{1}{3}=you\ can\ calculate\]